OpenStudy (anonymous):

Simplify 5y - (9y^2 + 5y - 6) +6y^2 I know you have to distribute right? But I don't know how to distribute when there is a 5y and - in front of the ( )

OpenStudy (anonymous):

no distribution.....just open the bracket by changing signs of all terms inside the bracket

OpenStudy (anonymous):

How do I change the signs of all terms?

OpenStudy (anonymous):

when there is a minus sign outside a bracket, you do as follows.. 5y - (9y^2 + 5y - 6) +6y^2 = 5y - 9y^2 - 5y + 6 + 6y^2 = - 9y^2 + 6y^2 - 5y + 5y + 6 = -3y^2 + 6

OpenStudy (anonymous):

Okay I'm gonna work it out and see if I get that!

OpenStudy (anonymous):

You will see that there are three terms inside the bracket and they change sign as follows: 9y^2 becomes - 9y^2 + 5y becomes - 5y and - 6 becomes + 6

OpenStudy (anonymous):

I don't understand why 5y is becoming negative

OpenStudy (anonymous):

because it was + 5y inside the bracket and on changing sign it becomes - 5y

OpenStudy (anonymous):

Okay so everything has to change signs when there is a negative in front of the brackets?

OpenStudy (anonymous):

correct, all the terms inside the brackets have to change sign

OpenStudy (anonymous):

But 6y^2 is still positive right?

OpenStudy (anonymous):

yeah, because it is OUTSIDE the bracket......

OpenStudy (anonymous):

Oh right! Sorry

OpenStudy (anonymous):

I'm gonna try to work it out now. Sorry my teacher never went over this and then put it on the quiz

OpenStudy (anonymous):

no problem, if u want I will give you a similar sum and you can solve and I will check if it is correct or not..

OpenStudy (anonymous):

Sure that sounds good! And would the answer to this problem be -9y^2 + 6y^2 +6?

OpenStudy (anonymous):

No, answer is -3y^2 + 6 because you have two like terms having y^2 so you have to simplify it.....

OpenStudy (anonymous):

Okay that's what I was thinking but I wasn't sure if they had to be simplified

OpenStudy (anonymous):

I understand now!

OpenStudy (anonymous):

solve this... -3p^4 + 5p^2 - (-2p^4 + 7p^3 + 2p^2) - 11p - (-5p + 2p^3 + 20) + 30

OpenStudy (anonymous):

Wow! Okay this going to take a while!

OpenStudy (anonymous):

Wow that was really hard I'm sure I got it wrong I got 3p^2 - 5p^4 - 9p^3 - 16p +10

OpenStudy (anonymous):

-3p^4 + 5p^2 - (-2p^4 + 7p^3 + 2p^2) - 11p - (-5p + 2p^3 + 20) + 30 = -3p^4 + 5p^2 + 2p^4 - 7p^3 - 2p^2 - 11p + 5p - 2p^3 - 20 + 30 = -3p^4 + 2p^4 - 7p^3 - 2p^3 + 5p^2 - 2p^2 - 11p + 5p - 20 + 30 = -p^4 - 9p^3 + 3p^2 - 6p + 10

OpenStudy (anonymous):

I don't get why it's +2p^4. I switched the signs to make 2 positives but then there was still the negative sign outside of the brackets so I kept it there

OpenStudy (anonymous):

the minus sign outside is "used" to change the signs of the terms inside and so it goes away....you just write the terms with changed signs as i hv done...

OpenStudy (anonymous):

Oh okay. And does it matter which order you put the terms in? Like for instance did you have to start with -3p^4 because it was the first one? Because in the other problem the 5y was first but then it moved to be after the 9y^2. So I'm just confused with which order to go in

OpenStudy (anonymous):

We write them in the decreasing order of their powers by rearranging the terms....peruse my solutions to both questions carefully....

OpenStudy (anonymous):

Oh okay got it! that makes sense. Thanks so mcuh! You were so helpful! And that was really nice of you to give me another problem

OpenStudy (anonymous):

I aim to "Teach" while helping........just giving solutions is the worst thing.......

OpenStudy (anonymous):

Yeah definitely! Thanks again!!!

OpenStudy (anonymous):

just a suggestion....always click "Best Response" button for those whose help u appreciate....